OpenStudy (anonymous):

If F(x) = x² + 5x and G(x) = 2x + 1, find F(5) + G(6)... please! :))))))))

OpenStudy (anonymous):

I just need an answer don't have to show work

jigglypuff314 (jigglypuff314):

Hello and Welcome to Openstudy this site does not allow the giving of answers, sorry :) read the code of conduct for more reference http://openstudy.com/code-of-conduct but I'll be more than willing to help! :D f(5) means that you plug in 5 for x into f(x) = x² + 5x so f(5) = (5)² + 5(5) = ? g(6) means that you plug in 6 for x into g(x) = 2x + 1 so g(6) = 2(6) + 1 = ? then once you get each, you add the two parts :)

OpenStudy (anonymous):

thank lots

jigglypuff314 (jigglypuff314):

glad I could help :)